OpenStudy (omnomnom):

The acceleration of gravity on the moon is about 1.6 m/sec 2. An experiment to test gravity compares the time it takes objects to reach a speed of 10 m/sec after being dropped from rest. How long does an object dropped on the moon have to fall to reach this speed compared to an object dropped on Earth? Select one: a. 1.6 times as long b. 6.1 times as long c. 9.8 times as long d. 15.7 times as long

OpenStudy (anonymous):

Okay so for the moon it is \(1.6m/s^2\) and on the earth it is \(9.81m/s^2\)

OpenStudy (omnomnom):

yeah

OpenStudy (anonymous):

But we want to find the ratio in terms of time.

OpenStudy (anonymous):

So first we have to find out how long it takes to fall on moon and how long it takes to fall on earth.

OpenStudy (omnomnom):

oh so it obliviously falls slower

OpenStudy (anonymous):

Yeah

OpenStudy (omnomnom):

so the logical answer would be a

OpenStudy (anonymous):

No, all of the answers assume it falls slower.

OpenStudy (anonymous):

We want to know how much slower

OpenStudy (omnomnom):

ohh we use an equation to solve it

OpenStudy (anonymous):

If we know \(v_i=0m/s, v_f=1.6m/s,a=9.81m/s^2\) how long will it take to fall on earth?

OpenStudy (omnomnom):

umm.....so we find time hold on

OpenStudy (anonymous):

\[\Delta v=v_f-v_i=at\]

OpenStudy (omnomnom):

its 0.16 seconds

OpenStudy (anonymous):

I made a mistake... \(v_f = 10m/s\)

OpenStudy (omnomnom):

Oh okay so it would be 1 second

OpenStudy (anonymous):

There are two \(a\)...\[\Large a_{earth}=9.81m/s^2, \ a_{moon} = 1.6m/s^2\]

OpenStudy (anonymous):

So now find the time it would take on the moon.

OpenStudy (omnomnom):

okay it would be what i said before 0.16 seconds

OpenStudy (omnomnom):

So the answer is a?

OpenStudy (anonymous):

No

OpenStudy (omnomnom):

why?

OpenStudy (anonymous):

Because 10/1.6 is not 0.16

OpenStudy (omnomnom):

oh so what do we do?

OpenStudy (anonymous):

We divide to get the correct time for the moon.

OpenStudy (omnomnom):

divide what? 10/1.6

OpenStudy (omnomnom):

if its that i got 6.25

OpenStudy (omnomnom):

so the answer is b?

OpenStudy (anonymous):

sorry I was gone. b is correct

OpenStudy (omnomnom):

i know :D